Section 18  Teaching and Popularization of Mathematics

Section 18: Teaching and Popularization of Mathematics

Invited Lectures

Monday, 24 August  MA 004

17:15 – 18:00 Mogens Niss, Roskilde University, Roskilde, Denmark Aspects of the nature and the state of research in mathematical education (I: p.198) Chair: Urs Kirchgraber

Tuesday, 25 August  H 104

14:00 – 14:45 Mariolina Bartolini Bussi, Università di Modena, Modena, Italy Drawing instruments: the theory of a practice from history to didactics (I: p.197) Chair: Klaus-D. Graf

15:00 – 15:45 James Stigler, University of California, Los Angeles, USA Classroom Mathematics Instruction in Three Cultures: An Introduction to the TIMSS Video Study (I: p.200) Chair: Thomas Jahnke

16:15 – 18:00 George Eyre Andrews, Pennsylvania State University, University Park, USA D.J. Lewis (moderator), NSF, Arlington, USA David Alexander Smith, Duke University, Durham, USA Collegiate Mathematics Instruction for the 21st Century — A Panel Discussion (I: p.200) Chair:

Wednesday, 26 August  H 104

14:00 – 15:45 Michèle Artigue (speaker), Université de Paris 7, Paris, France De la compréhension des processus d’appretissage à la conception de processus d’enseighement (I: p.202) Hans-Georg Steiner (respondent), Universität Bielefeld, Bielefeld, Germany “From the Understanding of Learning Processes to the Design of Teaching Processes” (I: p.203) Chair: Hans-Christian Reichel

16:15 – 18:00 Miguel de Guzman, Universidad Complutense de Madrid, Madrid, Spain Bernard R. Hodgson, Université Laval, Québec, Canada Aline Robert, IUFM de Versailles, France Vinicio Villani (moderator), Università di Pisa, Pisa, Italy Round table on: Difficulties in the Passage from Secondary to Tertiary Education (I: p.204) Chair:

Teaching and Popularization of Mathematics  Section 18

Short Communications and Poster Sessions

Short Communications  Saturday, 22 August
Undergraduate Mathematics at the University  H 2013

Chair: Joseph W. Eyles

16:15 – 16:30 Talma O. Leviatan, Tel Aviv University, Tel Aviv, Israel Misconceptions, fallacies, pitfalls, and paradoxes in the teaching of probability (p.361)

16:30 – 16:45 Lyudmila Y. Khilchenko, Yavneh Academy, Los Angeles, CA, USA An algorithmic approach to the notion of an elementary function (p.360)

16:45 – 17:00 Joseph W. Eyles, University of Texas at Austin, Austin, Texas, USA Discovering R.L. Moore’s Calculus Course (p.358)

Geometry in the Classroom  H 2013

Chair: F. Rudolf Fritsch

17:15 – 17:30 Klaus-D. Graf, Freie Universität Berlin, Berlin, Germany Distance learning fördert die Kreativität im Mathematikunterricht (p.360)

17:30 – 17:45 Bernd Zimmermann, Friedrich-Schiller-Universität Jena, Jena, Germany Al Sijzi on problemsolving and creativity in elementary geometry (p.364)

17:45 – 18:00 F. Rudolf Fritsch, Mathematisches Institut, München, Deutschland Remarks on quadrangles (p.359)

Short Communications  Monday, 24 August
Pedagogical and Psychological Aspects  H 2013

Chair: Hans-Joachim Arnold

14:00 – 14:15 Sou-yung Chiu, Chang-Hua University of Education, Chang-Hua, Taiwan Pedagogical Mathematics Knowledge, PMK and P(PMK)K (p.358)

14:15 – 14:30 Torsten Fritzlar, Friedrich-Schiller-Universität Jena, Jena, Germany Über Möglichkeiten der Sensibilisierung für die Komplexität von Mathematikunterricht (p.359)

14:30 – 14:45 Hans-Joachim Arnold, Universität Duisburg, Duisburg, Germany Geometrische Relationen-Algebra und Piaget’sche Schematheorie (p.357)

Miscellaneous I  H 2013

Chair: Masao Mori

15:00 – 15:15 Salem A. Sahab, King Abdul-Aziz University, Jeddah, Saudi Arabia Mathematics basic skills for grades 1 to 3 (p.363)

15:15 – 15:30 Josef Nietzsch, Humboldt–Universität zu Berlin, Berlin, Germany The Study of Mathematics by High School Students at the University (p.363)

15:30 – 15:45 Masao Mori, Chuo University, Tokyo, Japan Beginners Confusion and Terminology in Mathematics (p.362)

Miscellaneous II  H 2013

Chair: Mary E. Morley

16:15 – 16:30 Ludwig Paditz, HTW Dresden (FH), Dresden, Germany Comparision of the statistic calculators TI-83, EL-9600, and CFX-9850G PLUS (p.363)

16:30 – 16:45 Mary E. Morley, Educational Testing Service, Princeton, United States of America Three Response Types for Broadening the Conception of Mathematical Problem Solving in Computer-Graded Tests (p.362)

16:45 – 17:00 Zhongling Xu, National Natural Science Foundation of China, Beijing, P.R. of China The Popularization of Mathematical Methods in China (p.364)

Poster Session  Saturday, 22 August 14:00 – 15:45  MA 144

2  ICM’98 – Program

Section 18  Teaching and Popularization of Mathematics

Gloria L.M. Alves, Instituto Militar de Engenharia, Rio de Janeiro, Brazil

Modernization of Basic Disciplines in Engineering Courses at IME (p.357) Gerhard Brecht, Fachhochschule Lippe, Itzehoe, Germany

Do Mathematical Standards Concern Mathematics? (p.357) Cecilia Crespo Crespo, Department d’Algebra i Geometria, Barcelona, Spain

Thinking about Geometry Teaching (p.358) Paul Fjelstad, Northfield, MN, USA

A Picture for Real Arithmetic (p.359) Peter Kortesi, University of Miskolc, Miskolc, Hungary

Junior Mathematical Congress — Challenge for Future Mathematicians (p.360) Elena Vladimirovna Kuragina, Mari State University, Yoshkar-Ola, Russia

Methodical support for a programming-methodical complex of study methods of optimization (p.361) Julia M. Moshnikova, Karelian State Pedagogical University, Petrozavodsk, Russia

Correspondence School of Mathematics in Karelia (p.362) Yasuhiro Ozaki, Hachinohe Institute of Technology, Hachinohe-shi Aomori, Japan

An Approach in Teaching by Means of the Personal Computer Software “Mathematica” (p.363)

ICM’98 – Program  3